OpenStudy (anonymous):

If y = (x + 3)2, then (-2x - 6)2 must equal which of the following? A. -4y2 B. -2y2 C. -4y D. 4y wHAT FORMULA DO I USE TO GET THE ANSWER?

OpenStudy (anonymous):

Does anyone know what formula I use to get the answer?

OpenStudy (anonymous):

You use math... \[y = (x+3)^2\] \[\sqrt(y) = x+3\] \[(-2x-6)^2 = (-2\sqrt(y))^2\] =4y (...gotta learn how to write latex again)

OpenStudy (anonymous):

@nphuongsun93 Thank can you walk me through the steps so I know how to do it with other problems?

OpenStudy (anonymous):

Ok. \[y = (x+3)^2\]\[\sqrt(y) = x+3\] Understand this so far?

OpenStudy (jdoe0001):

hint: \( y={\color{brown}{ (x+3)^2 }}\qquad (-2x-6)^2\to \begin{cases} [-2(x+3)]^2\impliedby \textit{common factor}\\ (-2)^2(x+3)^2\impliedby expanded\\ +4{\color{brown}{ (x+3)^2 }} \end{cases}\)